OpenStudy (anonymous):

how do we integrate sin(x^2)dx?

OpenStudy (anonymous):

do by parts

OpenStudy (anonymous):

that might help !! :)

OpenStudy (amistre64):

try a sub of say: u = x^2; du = 2x dx; du/2sqrt(u) = dx maybe?

OpenStudy (amistre64):

might not work :)

OpenStudy (anonymous):

no parts doesnt help even a bit

OpenStudy (anonymous):

i tried ive been working on it for two days now

OpenStudy (amistre64):

http://www.wolframalpha.com/input/?i=integrate+sin%28x^2%29dx it takes some higher stuff

OpenStudy (anonymous):

whats that "s" thing?

OpenStudy (anonymous):

no trick is different

OpenStudy (anonymous):

ooops i should read first. it is \[\int \sin(x^2)dx\]?

OpenStudy (anonymous):

or \[\int \sin^2(x)dx\]?

OpenStudy (anonymous):

you will work on it for more than two days because you cannot do it.

OpenStudy (amistre64):

the "s" thing is a part of a higher course that I have yet to play in

OpenStudy (anonymous):

that S thing is just an integral. it is saying that your integral is an integral! so forget it.